OpenStudy (anonymous):

A student is asked to answer 8 out of 11 questions on an exam. (a) Classify the problem as a permutation, a combination, or neither. Combination Neither Permutation (b) In how many ways can she select the questions to be answered?

OpenStudy (amistre64):

there doesnt seem to be more than one group of options involved and it appears to be selection without repetition... I am guessing since its been awhile, but permutation. Would you happen to know the definitions of the words involved?

OpenStudy (amistre64):

permutation is is when the order that is picked matters. Combination tends to refer to without order. So this I think is actually leaning towards combination

OpenStudy (anonymous):

premutation is a selection of objects from a given set with regard to the order in which they are selected

OpenStudy (amistre64):

so the order to the questions selected is of no importance right?

OpenStudy (anonymous):

i guess so

OpenStudy (amistre64):

if you are given 10 answers to fill out and you can choose to do any 5 of them, does it matter what order you do it in?

OpenStudy (amistre64):

Id go with combination after reading up on the definitions :)

OpenStudy (amistre64):

we can write the rest of it in notation: 11! 9*10*11 ----- = -------- 8! 3! 1*2*3

OpenStudy (anonymous):

ok, do you possibly know how to do the second half?

OpenStudy (amistre64):

or 3*5*11 = 15*11 = 165

OpenStudy (amistre64):

let me check that again :)

OpenStudy (amistre64):

I wonder if id be different this way: 11! 4.5.6.7.8.9.10.11 ---- = ------------------ = same :) 3! 8! 1.2.3.4.5.6.7.8

OpenStudy (amistre64):

If I did it right, and thats a big if these days, the answers 165 ways

OpenStudy (anonymous):

ok, thank you for your help.